Make New Entry,
Make Followup Entry
User name jroche
Log entry time 08:04:31 on October 15, 2010
Entry number 334985
keyword=Owl shift summary
October 15, 2010, Owl shift summary
Taking calibration data on the dvcs calorimeter.
- 1:00 am: start shift with calorimeter elastic calibration in
progress. Pnom=2.406 GeV, thenom=37.00 deg. Beam current is 2.9 uA (per MCC value) .
Beam is centered A(x,y)=0,-1.3 and
B(x,y)=0,0. No raster. Beam energy is 5551.76 MeV. Ps5=1 all others
at 0. T5 rate is 9 Hz at 3 uA beam current.
- 12:00: Alex, Carlos and Magali are working on the DVCS DAQ.
- 00:30: taking one run in nominal condition at 4.7 uA: run 7379.
- 00:44: back at 3 uA.
- 2:00: Florian starts the checklist
- 2:20: MCC announces that they will do a bit of steering in the
- 3:00 time to change configuration: request access to check the
ladders are unlock. They are.
- 4:00 Trouble moving the spectro to the correct angle:
- New LHRS configuration is : Pnom=2.607 GeV, thenom=34.76
deg. T5 rate is 7.7 Hz at 3 uA beam current.
- 6:05. MCC looses beam and cannot restore it. They call their
experts and report that it's going to be a while... We are missing
20 min of data at this setting.
- 6:45: beam is back
- 7:30: trying to go to the next kinematic. The LHRS is stuck
again, paging Howard Smith. Howard says he is on is way.
- Elastic calibration (Pnom=2.406 GeV, thenom=37.00 deg, HV=500 V):
7381 (no ARS data :(, not sure why ) , 7382 , 7383
- Elastic calibration (Pnom=2.607 GeV, thenom=34.76 deg, HV=500 V):
7384, 7385, 7386